Ben invested $800, part at 9% per annum and the rest at 12% per annum. After one year the total interest earned was $79.50. How much did he invest?

we are solving them by elimination (gr 11)
i dont know how to do this one at all.

Answer by mananth(16949)

Ben invested $800, part at 9% per annum and the rest at 12% per annum. After one year the total interest earned was $79.50. How much did he invest?

9% per annum ------ x
12% per annum--------y
x+y =800....1
0.09x+0.12y=$79.50.
multiply by 100
9x+12y=7950............(2)
multiply (1)--9
-9x-9y=-7200
add to (2)
+3y=750
/3
y = $250 ===> at 12%
Balance $550 ay 9%